# Learning hidden cascades via classification

### v2

@@ Appendix B Algorithms
2:Input: $\text{label}\in\{0,1\}$ indicates observed (00) and simulated($1$) data
[…]
@@
7:         for each simulation $n=1$ to $N$ do $\triangleright$ $N$: Number of cascades
[…]
@@
13:              Extract symptom vector $\mathbf{s}_{n}\in\mathbb{R}^{E}$triangleright $E$: Number of entities
@@
[…]
@@
3:$\mathcal{F}_{\text{GT}}\leftarrow$ GenerateFeatureVector(label(0), $\theta$) $\triangleright$ Ground truth features
@@
7:         $\text{accuracy}[v]\leftarrow f^{(v)}_{\text{classifier}}(\mathcal{F}_{\text{GT}}[v],\mathcal{F}_{\text{pred}}[v])$triangleright Entity-specific classifier
@@
5:     $\psi_{v}\leftarrow$ FindHyperparam($f^{(v)}_{\text{classifier}}(\mathcal{F}_{\text{GT}}[v],\mathcal{F}_{\text{pred}}[v])$) $\triangleright$ Tune classifier hyperparameters
@@
1:Input: $\mathcal{C}=\{c_{1},c_{2},\dots,c_{28}\}$triangleright Set of companies
@@
3:Output: $\mathcal{F}^{\text{train}}_{\text{GT}},\mathcal{F}^{\text{test}}_{\text{GT}}$triangleright Training and testing feature vectors
@@
10:              $\mathcal{T}_{i}\leftarrow\mathcal{A}_{i}$, $\Delta\leftarrow 1$triangleright Announcement days and 1-day return
@@
12:              $\mathcal{T}_{i}\leftarrow$ non-announcement trade days, $\Delta\leftarrow 5$triangleright Regular days and 5-day return
@@
15:              Initialize vector $\mathbf{v}_{t}\leftarrow[]$triangleright Vector for all investors at time $t$
@@
24:                       $\mathbf{v}_{t}\leftarrow\mathbf{v}_{t}\cup\left\{\begin{array}[]{ll}+1,&\text{if }p_{u}<p_{t}^{+\Delta}\\ -1,&\text{otherwise}\end{array}\right.$triangleright Profitability encoding
@@
26:                       $\mathbf{v}_{t}\leftarrow\mathbf{v}_{t}\cup\{0\}$triangleright No trade found
@@
Appendix C Hyperparameter Search Space
@@
[…]
@@
Appendix D Network Structures
@@
Appendix E Tables
Tables 2, 3, and 4 report results from simulation experiments conducted on synthetic networks with known ground truth. These experiments vary the propagation probability $p$ and symptom (or observation) probability $q$, and are executed on different graph structures: tree, loopy, and empirical graph topologies. Importantly, these simulations focus purely on the structural properties of the networks and do not incorporate investor behavior or transaction data.
@@ Appendix E Tables
Table 2: Robustness Check for Tree Across Multiple Values of $p$ and $q$
$q$ (0.1)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.13 ± 0.040 0.10 ± 0.000 $1.20\times 10^{-3}$ 0.43%
0.3 0.29 ± 0.020 0.12 ± 0.020 $7.03\times 10^{-4}$ 0.43%
$p$ 0.5 0.46 ± 0.060 0.10 ± 0.000 $2.22\times 10^{-3}$ 0.43%
0.7 0.70 ± 0.005 0.10 ± 0.000 $1.62\times 10^{-5}$ 0.43%
0.9 0.90 ± 0.004 0.10 ± 0.009 $6.44\times 10^{-5}$ 0.43%
$q$ (0.3)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.15 ± 0.030 0.36 ± 0.140 $1.17\times 10^{-2}$ 0.43%
0.28 ± 0.030 0.31 ± 0.020 $7.62\times 10^{-4}$ 0.43%
0.49 ± 0.007 0.29 ± 0.020 $2.94\times 10^{-4}$ 0.43%
0.70 ± 0.004 0.30 ± 0.005 $3.64\times 10^{-5}$ 0.43%
0.90 ± 0.000 0.30 ± 0.000 $5.14\times 10^{-6}$ 0.43%
$q$ (0.5)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.12 ± 0.023 0.54 ± 0.069 $3.11\times 10^{-3}$ 0.43%
0.3 0.29 ± 0.030 0.51 ± 0.036 $1.05\times 10^{-3}$ 0.43%
$p$ 0.5 0.44 ± 0.044 0.57 ± 0.112 $9.64\times 10^{-3}$ 0.44%
0.7 0.68 ± 0.040 0.55 ± 0.112 $6.67\times 10^{-3}$ 0.44%
0.9 0.90 ± 0.006 0.50 ± 0.000 $7.73\times 10^{-6}$ 0.43%
$q$ (0.7)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.14 ± 0.058 0.68 ± 0.051 $3.37\times 10^{-3}$ 0.43%
0.29 ± 0.055 0.74 ± 0.096 $5.58\times 10^{-3}$ 0.43%
0.49 ± 0.016 0.69 ± 0.024 $3.70\times 10^{-4}$ 0.43%
0.69 ± 0.019 0.73 ± 0.055 $1.76\times 10^{-3}$ 0.43%
0.88 ± 0.030 0.74 ± 0.077 $3.10\times 10^{-3}$ 0.47%
$q$ (0.9)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.12 ± 0.040 0.73 ± 0.080 $1.71\times 10^{-2}$ 0.43%
0.3 0.30 ± 0.010 0.90 ± 0.010 $9.44\times 10^{-5}$ 0.43%
$p$ 0.5 0.51 ± 0.020 0.90 ± 0.070 $2.08\times 10^{-3}$ 0.43%
0.7 0.68 ± 0.030 0.92 ± 0.040 $1.18\times 10^{-3}$ 0.49%
0.9 0.95 ± 0.050 0.84 ± 0.080 $7.45\times 10^{-3}$ 0.74%
$q$ (0.1)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.10 ± 0.005 0.10 ± 0.004 $1.98\times 10^{-5}$ 0.43%
$p$ 0.3 0.30 ± 0.000 0.10 ± 0.005 $1.24\times 10^{-5}$ 0.43%
0.5 0.50 ± 0.004 0.10 ± 0.000 $1.23\times 10^{-5}$ 0.44%
0.7 0.70 ± 0.005 0.10 ± 0.000 $2.05\times 10^{-5}$ 0.43%
0.9 0.90 ± 0.005 0.10 ± 0.000 $3.06\times 10^{-5}$ 0.43%
$q$ (0.3)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.017 0.28 ± 0.019 $4.30\times 10^{-4}$ 0.43%
0.30 ± 0.009 0.31 ± 0.005 $4.79\times 10^{-5}$ 0.43%
0.50 ± 0.006 0.30 ± 0.000 $1.40\times 10^{-5}$ 0.43%
0.70 ± 0.000 0.30 ± 0.000 $3.45\times 10^{-6}$ 0.43%
0.89 ± 0.014 0.30 ± 0.000 $6.73\times 10^{-5}$ 0.43%
$q$ (0.5)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.09 ± 0.015 0.53 ± 0.052 $1.78\times 10^{-3}$ 0.43%
0.3 0.30 ± 0.005 0.50 ± 0.017 $1.44\times 10^{-4}$ 0.48%
$p$ 0.5 0.50 ± 0.006 0.50 ± 0.000 $9.17\times 10^{-6}$ 0.43%
0.7 0.70 ± 0.000 0.50 ± 0.000 $6.93\times 10^{-6}$ 0.43%
0.9 0.90 ± 0.005 0.50 ± 0.000 $1.33\times 10^{-5}$ 0.43%
$q$ (0.7)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.023 0.70 ± 0.057 $1.69\times 10^{-3}$ 0.43%
0.30 ± 0.000 0.70 ± 0.010 $7.91\times 10^{-5}$ 0.42%
0.50 ± 0.000 0.70 ± 0.000 $4.06\times 10^{-6}$ 0.43%
0.70 ± 0.000 0.70 ± 0.000 $2.37\times 10^{-6}$ 0.43%
0.90 ± 0.007 0.70 ± 0.000 $2.27\times 10^{-5}$ 0.43%
$q$ (0.9)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.11 ± 0.021 0.84 ± 0.109 $6.74\times 10^{-3}$ 0.44%
0.3 0.31 ± 0.005 0.87 ± 0.025 $9.10\times 10^{-4}$ 0.43%
$p$ 0.5 0.50 ± 0.000 0.90 ± 0.000 $2.76\times 10^{-7}$ 0.43%
0.7 0.70 ± 0.000 0.90 ± 0.000 $1.35\times 10^{-6}$ 0.43%
0.9 0.91 ± 0.008 0.90 ± 0.000 $4.50\times 10^{-5}$ 0.43%
$q$ (0.1)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.10 ± 0.009 0.10 ± 0.005 $7.43\times 10^{-5}$ $1.62\times 10^{-2}$%
0.3 0.37 ± 0.027 0.10 ± 0.009 $2.66\times 10^{-3}$ $1.07\times 10^{-2}$%
$p$ 0.5 0.54 ± 0.024 0.10 ± 0.005 $8.67\times 10^{-4}$ $6.04\times 10^{-3}$%
0.7 0.71 ± 0.007 0.10 ± 0.008 $1.07\times 10^{-4}$ $4.83\times 10^{-3}$%
0.9 0.89 ± 0.022 0.10 ± 0.005 $2.21\times 10^{-4}$ $4.75\times 10^{-3}$%
$q$ (0.3)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.012 0.30 ± 0.009 $1.35\times 10^{-4}$ $1.66\times 10^{-2}$%
0.38 ± 0.041 0.30 ± 0.010 $3.74\times 10^{-3}$ $1.16\times 10^{-2}$%
0.55 ± 0.059 0.30 ± 0.007 $2.69\times 10^{-3}$ $6.96\times 10^{-3}$%
0.72 ± 0.027 0.30 ± 0.008 $5.98\times 10^{-4}$ $6.45\times 10^{-3}$%
0.91 ± 0.025 0.30 ± 0.011 $3.45\times 10^{-4}$ $5.88\times 10^{-3}$%
$q$ (0.5)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.11 ± 0.013 0.49 ± 0.008 $1.16\times 10^{-4}$ $1.68\times 10^{-2}$%
0.3 0.36 ± 0.033 0.50 ± 0.004 $2.43\times 10^{-3}$ $1.18\times 10^{-2}$%
$p$ 0.5 0.56 ± 0.036 0.50 ± 0.005 $2.36\times 10^{-3}$ $7.89\times 10^{-3}$%
0.7 0.75 ± 0.019 0.50 ± 0.011 $1.34\times 10^{-3}$ $7.08\times 10^{-3}$%
0.9 0.91 ± 0.036 0.50 ± 0.000 $6.06\times 10^{-4}$ $6.57\times 10^{-3}$%
$q$ (0.7)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.11 ± 0.004 0.69 ± 0.016 $2.80\times 10^{-4}$ $1.67\times 10^{-2}$%
0.37 ± 0.019 0.69 ± 0.005 $2.53\times 10^{-3}$ $1.18\times 10^{-2}$%
0.53 ± 0.021 0.70 ± 0.007 $7.28\times 10^{-4}$ $7.12\times 10^{-3}$%
0.72 ± 0.027 0.70 ± 0.015 $5.41\times 10^{-4}$ $6.71\times 10^{-3}$%
0.91 ± 0.046 0.70 ± 0.005 $8.15\times 10^{-4}$ $6.34\times 10^{-3}$%
$q$ (0.9)
$\hat{p}$ $\hat{q}$ $\overline{MSE}$ $\overline{CA}$
0.1 0.11 ± 0.005 0.90 ± 0.011 $7.49\times 10^{-5}$ $1.65\times 10^{-2}$%
0.3 0.33 ± 0.013 0.90 ± 0.008 $5.44\times 10^{-4}$ $1.06\times 10^{-2}$%
p 0.5 0.51 ± 0.004 0.90 ± 0.005 $6.42\times 10^{-5}$ $5.89\times 10^{-3}$%
0.7 0.72 ± 0.019 0.91 ± 0.005 $5.01\times 10^{-4}$ $5.44\times 10^{-3}$%
0.9 0.91 ± 0.022 0.91 ± 0.005 $2.74\times 10^{-4}$ $4.70\times 10^{-3}$%
@@
Company 4 $3.93\times 10^{-3}$ $3.34\times 10^{-3}$ $9.93\times 10^{-1}$
@@
Company 1 0.39 0.12 $2.22\times 10^{-2}$0.18 0.02 $1.19\times 10^{-2}$2.09 6.30
Company 2 0.59 0.03 $2.93\times 10^{-2}$0.30 0.01 $1.41\times 10^{-2}$1.97 2.07
Company 3 0.60 0.04 $2.57\times 10^{-2}$0.32 0.02 $1.89\times 10^{-2}$1.86 1.69
Company 4 0.53 0.03 $3.30\times 10^{-2}$0.29 0.06 $1.41\times 10^{-2}$1.81 0.53
Company 5 0.40 0.06 $1.85\times 10^{-2}$0.23 0.05 $1.21\times 10^{-2}$1.74 1.08
Company 6 0.50 0.03 $2.04\times 10^{-2}$0.31 0.02 $1.72\times 10^{-2}$1.63 1.60
Company 7 0.57 0.03 $2.21\times 10^{-2}$0.36 0.04 $2.58\times 10^{-2}$1.61 0.66
Company 8 0.52 0.03 $2.64\times 10^{-2}$0.35 0.06 $2.67\times 10^{-2}$1.48 0.45
Company 9 0.57 0.02 $1.96\times 10^{-2}$0.38 0.03 $2.32\times 10^{-2}$1.47 0.73
Company 10 0.47 0.04 $2.71\times 10^{-2}$0.36 0.03 $1.64\times 10^{-2}$1.30 1.14
Company 11 0.35 0.09 $2.00\times 10^{-2}$0.28 0.05 $1.52\times 10^{-2}$1.26 1.79
Company 12 0.36 0.03 $1.94\times 10^{-2}$0.29 0.01 $1.76\times 10^{-2}$1.23 2.08
Company 13 0.58 0.05 $3.56\times 10^{-2}$0.48 0.02 $3.15\times 10^{-2}$1.22 2.63
Company 14 0.55 0.04 $1.91\times 10^{-2}$0.46 0.02 $1.59\times 10^{-2}$1.22 2.57
Company 15 0.49 0.02 $1.66\times 10^{-2}$0.41 0.02 $2.20\times 10^{-2}$1.19 1.37
Company 16 0.35 0.06 $2.20\times 10^{-2}$0.31 0.04 $1.48\times 10^{-2}$1.13 1.40
Company 17 0.36 0.04 $2.75\times 10^{-2}$0.33 0.01 $1.28\times 10^{-2}$1.07 2.73
Company 18 0.42 0.05 $9.61\times 10^{-3}$0.40 0.06 $2.30\times 10^{-2}$1.07 0.76
Company 19 0.36 0.07 $2.33\times 10^{-2}$0.35 0.03 $2.07\times 10^{-2}$1.04 2.62
Company 20 0.48 0.03 $2.44\times 10^{-2}$0.48 0.02 $2.49\times 10^{-2}$1.01 1.69
Company 21 0.41 0.03 $1.87\times 10^{-2}$0.41 0.01 $1.57\times 10^{-2}$1.00 2.32
Company 22 0.44 0.04 $3.49\times 10^{-2}$0.45 0.01 $1.85\times 10^{-2}$0.99 2.80
Company 23 0.42 0.06 $4.04\times 10^{-2}$0.44 0.01 $1.92\times 10^{-2}$0.95 4.06
Company 24 0.42 0.03 $1.66\times 10^{-2}$0.45 0.05 $2.01\times 10^{-2}$0.94 0.55
Company 25 0.41 0.03 $2.64\times 10^{-2}$0.44 0.01 $1.86\times 10^{-2}$0.93 2.34
Company 26 0.34 0.04 $2.38\times 10^{-2}$0.38 0.01 $1.71\times 10^{-2}$0.89 3.04
Company 27 0.45 0.09 $2.48\times 10^{-2}$0.66 0.01 $1.76\times 10^{-2}$0.68 7.90
Company 28 0.34 0.04 $3.34\times 10^{-2}$0.52 0.03 $4.69\times 10^{-2}$0.65 1.17

### v3

@@ Appendix D Algorithms
2:Input: $\text{label}\in\{0,1\}$ indicates observed (0) and simulated(11) data
3:Output: $\{\text{feature\_vector}_{m}\}_{m=1}^{M}$: Set of $M$ labeled feature vectors
@@
7:   for each simulation $n=1$ to NN do $\triangleright$ NN: Number of cascades
8:     if $\text{label}==1$ then
@@
13:     Extract symptom vector $\mathbf{s}_{n}\in\mathbb{R}^{E}$ $\triangleright$ $E$: Number of entities
@@
17:   $\text{feature\_vector}_{m}\leftarrow\text{append}(\mathbf{f}_{m},\text{label})\in\mathbb{R}^{E\times(d+1)}$
@@
[…]
@@
7:   $\text{accuracy}[v]\leftarrow f^{(v)}_{\text{classifier}}(\mathcal{F}_{\text{GT}}[v],\mathcal{F}_{\text{pred}}[v])$ $\triangleright$ Entity-specific classifier
@@
[…]
@@
1:Input: $\mathcal{C}=\{c_{1},c_{2},\dots,c_{28}\}$ $\triangleright$ Set of companies
@@
3:Output: $\mathcal{F}^{\text{train}}_{\text{GT}},\mathcal{F}^{\text{test}}_{\text{GT}}$ $\triangleright$ Training and testing feature vectors
@@
10:     $\mathcal{T}_{i}\leftarrow\mathcal{A}_{i}$, $\Delta\leftarrow 1$ $\triangleright$ Announcement days and 1-day return
@@
12:     $\mathcal{T}_{i}\leftarrow$ non-announcement trade days, $\Delta\leftarrow 5$ $\triangleright$ Regular days and 5-day return
@@
15:     Initialize vector $\mathbf{v}_{t}\leftarrow[]$ $\triangleright$ Vector for all investors at time $t$
@@
24:        $\mathbf{v}_{t}\leftarrow\mathbf{v}_{t}\cup\left\{\begin{array}[]{ll}+1,&\text{if }p_{u}<p_{t}^{+\Delta}\\ -1,&\text{otherwise}\end{array}\right$. $\triangleright$ Profitability encoding
@@
26:        $\mathbf{v}_{t}\leftarrow\mathbf{v}_{t}\cup\{0\}$ $\triangleright$ No trade found
@@
Appendix E Hyperparameter Search Space
@@
  • $k$-Nearest Neighbors (KNN): $n\_neighbors\in\{5,10\},$ weights $\in\{\text{uniform},\text{distance}\},$ algorithm $\in\{\text{auto},\text{ball\_tree}\}$
@@
Appendix F Network Structures
@@
Appendix G Descriptive statistics and Results on Empirical Analysis
@@
Company 4 $93\times 10^{-3} 3.3$$34\times 10^{-3}$ $9.93\times 10^{-1}$
@@
Company 1 0.39 0.12 $2.22\times 10^{-2}$% 0.18 0.02 $1.19\times 10^{-2}$% 2.09 6.30
Company 2 0.59 0.03 $2.93\times 10^{-2}$% 0.30 0.01 $1.41\times 10^{-2}$% 1.97 2.07
Company 3 0.60 0.04 $2.57\times 10^{-2}$% 0.32 0.02 $1.89\times 10^{-2}$% 1.86 1.69
Company 4 0.53 0.03 $3.30\times 10^{-2}$% 0.29 0.06 $1.41\times 10^{-2}$% 1.81 0.53
Company 5 0.40 0.06 $1.85\times 10^{-2}$% 0.23 0.05 $1.21\times 10^{-2}$% 1.74 1.08
Company 6 0.50 0.03 $2.04\times 10^{-2}$% 0.31 0.02 $1.72\times 10^{-2}$% 1.63 1.60
Company 7 0.57 0.03 $2.21\times 10^{-2}$% 0.36 0.04 $2.58\times 10^{-2}$% 1.61 0.66
Company 8 0.52 0.03 $2.64\times 10^{-2}$% 0.35 0.06 $2.67\times 10^{-2}$% 1.48 0.45
Company 9 0.57 0.02 $1.96\times 10^{-2}$% 0.38 0.03 $2.32\times 10^{-2}$% 1.47 0.73
Company 10 0.47 0.04 $2.71\times 10^{-2}$% 0.36 0.03 $1.64\times 10^{-2}$% 1.30 1.14
Company 11 0.35 0.09 $2.00\times 10^{-2}$% 0.28 0.05 $1.52\times 10^{-2}$% 1.26 1.79
Company 12 0.36 0.03 $1.94\times 10^{-2}$% 0.29 0.01 $1.76\times 10^{-2}$% 1.23 2.08
Company 13 0.58 0.05 $3.56\times 10^{-2}$% 0.48 0.02 $3.15\times 10^{-2}$% 1.22 2.63
Company 14 0.55 0.04 $1.91\times 10^{-2}$% 0.46 0.02 $1.59\times 10^{-2}$% 1.22 2.57
Company 15 0.49 0.02 $1.66\times 10^{-2}$% 0.41 0.02 $2.20\times 10^{-2}$% 1.19 1.37
Company 16 0.35 0.06 $2.20\times 10^{-2}$% 0.31 0.04 $48\times 10^{-2}1.1$3 1.40
Company 17 0.36 0.04 $2.75\times 10^{-2}$% 0.33 0.01 $1.28\times 10^{-2}$% 1.07 2.73
Company 18 0.42 0.05 $9.61\times 10^{-3}$% 0.40 0.06 $2.30\times 10^{-2}$% 1.07 0.76
Company 19 0.36 0.07 $2.33\times 10^{-2}$% 0.35 0.03 $2.07\times 10^{-2}$% 1.04 2.62
Company 20 0.48 0.03 $2.44\times 10^{-2}$% 0.48 0.02 $2.49\times 10^{-2}$% 1.01 1.69
Company 21 0.41 0.03 $1.87\times 10^{-2}$% 0.41 0.01 $1.57\times 10^{-2}$% 1.00 2.32
Company 22 0.44 0.04 $3.49\times 10^{-2}$% 0.45 0.01 $1.85\times 10^{-2}$% 0.99 2.80
Company 23 0.42 0.06 $4\times 10^{-2}0.44 0$.01 $1.92\times 10^{-2}$% 0.95 4.06
Company 24 0.42 0.03 $1.66\times 10^{-2}$% 0.45 0.05 $2.01\times 10^{-2}$% 0.94 0.55
Company 25 0.41 0.03 $2.64\times 10^{-2}$% 0.44 0.01 $1.86\times 10^{-2}$% 0.93 2.34
Company 26 0.34 0.04 $2.38\times 10^{-2}$% 0.38 0.01 $1.71\times 10^{-2}$% 0.89 3.04
Company 27 0.45 0.09 $2.48\times 10^{-2}$% 0.66 0.01 $1.76\times 10^{-2}$% 0.68 7.90
Company 28 0.34 0.04 $3.34\times 10^{-2}$% 0.52 0.03 $4.69\times 10^{-2}$% 0.65 1.17